OpenStudy (anonymous):

Out of two thousand families with four children each, how many would you expect to have one or two girls? Assume the probability of a male birth is 1/2

OpenStudy (perl):

lets look at 1 family with 4 children

OpenStudy (perl):

the probability of having one or two children is P( X = 1 or X = 2) = P( X=1) + P(X=2)

OpenStudy (perl):